A cubical rice granary has an edge of 1.8 meters. How many sack of rice can it contain if one sack is 48 cubic decimeters?
Answer by CubeyThePenguin(3113)

Volume = (1.8 m)^3 = 5.832 m^3 = 5832 dm^3

5832/48 = 121.5

121 sacks (round down to the nearest integer)